### v1

# Deep Positive-Unlabeled Anomaly Detection for Contaminated Unlabeled Data

Abstract
Semi-supervised anomaly detection, which aims to improve the performance of the anomaly detector by using a small amount of anomaly data in addition to unlabeled data, has attracted attention. Existing semi-supervised approaches assume that unlabeled data are mostly normal. They train the anomaly detector to minimize the anomaly scores for the unlabeled data, and to maximize those for the anomaly data. However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the detection performance. To solve this problem, we propose the positive-unlabeled autoencoder, which is based on positive-unlabeled learning and the anomaly detector such as the autoencoder. With our approach, we can approximate the anomaly scores for normal data using the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the anomaly detector to minimize the anomaly scores for normal data, and to maximize those for the anomaly data. In addition, our approach is applicable to various anomaly detectors such as the DeepSVDD. Experiments on various datasets show that our approach achieves better detection performance than existing approaches.
1 Introduction
Anomaly detection, which aims to identify unusual data points, is an important task in machine learning [48]. It has been performed in various fields [24, 11, 28, 1, 26, 3, 30, 4, 42], such as cyber-security, fault and damage detection, novelty detection, fraud detection, and medical diagnosis.
In general, anomaly detection is performed by unsupervised learning because it does not require expensive and time-consuming labeling. Unsupervised anomaly detection assumes that unlabeled data are mostly normal, and tries to detect anomalies by using the difference from them. As an example, we focus on the autoencoder (AE) [16], which has been successfully applied to anomaly detection [50]. The AE learns the low-dimensional representation of data points through data reconstruction. The AE consists of two neural networks: the encoder and the decoder. The encoder maps a data point to the representation, and the decoder reconstructs the original data point from the representation. The AE is trained to minimize the reconstruction error, which is the difference between the data point and its reconstructed data point. After training, the AE is expected to successfully reconstruct normal data and fail to reconstruct anomaly data because the training dataset is assumed to contain only normal data and no anomaly data. Hence, the reconstruction error can be used for the anomaly score. Although unsupervised approaches are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies.
To improve the anomaly detection performance, semi-supervised anomaly detection uses a small amount of anomaly data in addition to unlabeled data. Existing approaches train the anomaly detector to minimize the anomaly scores for the unlabeled data, and to maximize those for the anomaly data [15, 47, 59]. For example, the autoencoding binary classifier (ABC) [59] trains the AE to minimize the reconstruction errors for the unlabeled data, and to maximize those for the anomaly data. However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the anomaly detection performance. This frequently occurs because it is difficult to label all anomalies.
Our purpose is to propose a semi-supervised learning approach that can improve the anomaly detection performance even if the unlabeled data are contaminated with anomalies. To handle such unlabeled data, we propose the positive-unlabeled autoencoder (PUAE), which is based on positive-unlabeled (PU) learning [9, 8, 22] and the anomaly detector such as the AE. PU learning assumes that an unlabeled data distribution is a mixture of normal and anomaly data distributions111 Note that the anomaly data in the training dataset follow the anomaly data distribution, but new types of anomalies, not included in the training dataset, do NOT follow this distribution. In general, no distribution can fully represent all possible anomalies. . Accordingly, the normal data distribution is approximated by the unlabeled and anomaly data distributions. With this assumption, we approximate the reconstruction errors for normal data with the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the AE to minimize the reconstruction errors for normal data, and to maximize those for the anomaly data.
(a) PU learning
(b) DAE
(c) ABC
(d) Proposed
Figure 1: The comparison of the PU learning, the unsupervised anomaly detector (DAE), the semi-supervised anomaly detector (ABC), and the proposed method on the toy dataset. The unlabeled data in this dataset include both normal and anomaly data points. The yellow and blue in the contour maps represent abnormality and normality, respectively. The purple stars represent the examples of unseen anomalies, which are new types of anomalies unseen during training.
Figure 1 shows the difference between the PU learning [22], the denoising AE (DAE) [55], the ABC, and the proposed method on the toy dataset. The DAE is a variant of the AE, which tries to reconstruct original data points from noisy input data points. This has also been successfully applied to anomaly detection [50]. Here, the ABC and the proposed method are based on the DAE. The toy dataset consists of the unlabeled and anomaly data, where the unlabeled data include both normal and anomaly data points. We first focus on the PU learning, which aims to train the binary classifier from the unlabeled and anomaly data. This can detect seen anomalies that are included in the training dataset. However, since its decision boundary is between normal data points and seen anomalies, it cannot detect unseen anomalies, which are new types of anomalies unseen during training, such as novel anomalies and zero-day attacks [56, 36, 7]. We next focus on the DAE and the ABC. They can detect unseen anomalies to some extent since they try to detect anomalies by using the difference from normal data. The DAE cannot detect seen anomalies because it cannot use information about anomalies. The ABC can detect seen anomalies to some extent since it can use the anomaly data. However, the contaminated dataset weakens the effect of maximizing the anomaly scores for anomalies in the ABC. Finally, we focus on the proposed method. The proposed method can detect seen anomalies according to the effectiveness of PU learning, and can detect unseen anomalies to some extent according to the effectiveness of the DAE.
Furthermore, our approach is applicable to various anomaly detectors. When choosing the detector, our approach requires that its loss function be non-negative and differentiable. A lot of approaches satisfy this requirement, such as the deep support vector data description (DeepSVDD) [46]. Our approaches achieve better detection performance than existing approaches on various datasets.
Our contributions can be summarized as follows:
  • We propose the positive-unlabeled autoencoder (PUAE), the semi-supervised anomaly detection approach that can handle contaminated unlabeled data.
  • Our approach is applicable to various anomaly detectors such as the DeepSVDD.
  • Experiments on various datasets show that our approach achieves better anomaly detection performance than existing approaches.
2 Preliminaries
In this section, we first explain our problem setup. Next, we review the autoencoder (AE) [16] and the autoencoding binary classifier (ABC) [59]. They are typical unsupervised and semi-supervised anomaly detection approaches, and our approach is based on them.
2.1 Problem Setup
Given unlabeled dataset $\mathcal{U}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ and anomaly dataset $\mathcal{A}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{M}\}$ for training. $\mathcal{U}$ contains not only normal data points but also seen anomalies included in $\mathcal{A}$. The test dataset contains normal data points, seen anomalies, and unseen anomalies not included in $\mathcal{A}$. Our goal is to obtain a high performance anomaly detector using $\mathcal{U}$ and $\mathcal{A}$.
2.2 Autoencoder
For unsupervised anomaly detectors, we train them only using the dataset $\mathcal{U}$. As an example, we focus on the AE, which has been successfully applied to anomaly detection [50]. The AE is presented for representation learning, which learns the representation of data points through data reconstruction. Let $\mathbf{x}$ be a data point and $\mathbf{z}$ be its low-dimensional latent representation. The AE consists of two neural networks: encoder $E_{\theta}(\mathbf{x})$ and decoder $D_{\theta}(\mathbf{z})$, where $\theta$ is the parameter of these neural networks. $E_{\theta}(\mathbf{x})$ maps a data point $\mathbf{x}$ into a low-dimensional latent representation $\mathbf{z}$, and $D_{\theta}(\mathbf{z})$ reconstructs the original data point $\mathbf{x}$ from the latent representation $\mathbf{z}$. The reconstruction error for each data point $\mathbf{x}$ in the AE is defined as follows:
$\ell(\mathbf{x};\theta)=\left\|D_{\theta}(E_{\theta}(\mathbf{x}))-\mathbf{x}\right\|,$ (1)
[…]
When the AE is used for anomaly detection, all unlabeled data points are assumed to be normal. We train the AE by minimizing the following objective function:
$\mathcal{L}_{\mathrm{AE}}(\theta)=\frac{1}{N}\sum_{n=1}^{N}\ell(\mathbf{x}_{n};\theta).$ (2)
After training, the AE is expected to successfully reconstruct normal data and fail to reconstruct anomaly data because the training dataset is assumed to contain only normal data and no anomaly data. Hence, the reconstruction error can be used for the anomaly score.
Although unsupervised approaches are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies.
2.3 Autoencoding Binary Classifier
Semi-supervised anomaly detection aims to improve the anomaly detection performance by using the datasets $\mathcal{U}$ and $\mathcal{A}$. A number of studies have been presented such as the ABC [59], the deep semi-supervised anomaly detection (DeepSAD) [47], and the outlier exposure [15]. Here, we focus on the ABC, which is based on the AE.
Let $y=0$ be normal and $y=1$ be anomaly. The ABC models the conditional probability of the normality or the abnormality $y$ given the data point $\mathbf{x}$ by using the reconstruction error $\ell(\mathbf{x};\theta)$ as follows:
$p_{\theta}(y|\mathbf{x})$ =$=\begin{cases}\exp(-\ell(\mathbf{x};\theta))&(y=0)\\ 1-\exp(-\ell(\mathbf{x};\theta))&(y=1)\end{cases}.$ (3)
A small reconstruction error results in a higher $p_{\theta}(y=0|\mathbf{x})$, while a large reconstruction error results in a higher $p_{\theta}(y=1|\mathbf{x})$.
With this conditional probability, the ABC introduces the binary cross entropy as the loss function for each data point as follows:
$\ell_{\mathrm{BCE}}(\mathbf{x},y;\theta)=-\log p_{\theta}(y|\mathbf{x})=(1-y)\ell(\mathbf{x};\theta)-y\log(1-\exp(-\ell(\mathbf{x};\theta))).$ (4)
Like the AE, the ABC assumes all unlabeled data points to be normal. We train the ABC by minimizing the following objective function:
$\mathcal{L}_{\mathrm{ABC}}(\theta)=\frac{1}{N}\sum_{n=1}^{N}\ell_{\mathrm{BCE}}(\mathbf{x}_{n},0;\theta)+\frac{1}{M}\sum_{m=1}^{M}\ell_{\mathrm{BCE}}(\tilde{\mathbf{x}}_{m},1;\theta).$ (5)
This minimizes the reconstruction errors for the unlabeled data and maximizes those for the anomaly data. Hence, after training, the AE becomes to reconstruct the unlabeled data assumed to be normal, and fail to reconstruct anomalies. Other semi-supervised anomaly detection approaches such as the DeepSAD [47] and the outlier exposure [15] also minimize the anomaly scores for the unlabeled data and maximize those for the anomaly data.
However, the unlabeled dataset $\mathcal{U}$ is often contaminated with anomalies in practice. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the anomaly detection performance. This frequently occurs because it is difficult to label all anomalies.
3 Proposed Method
We aim to improve the detection performance even if the unlabeled dataset $\mathcal{U}$ contains anomalies. To handle such unlabeled dataset, we propose the positive-unlabeled autoencoder (PUAE), which is based on positive-unlabeled (PU) learning [9, 8, 22] and the anomaly detector such as the AE. Hereinafter, we also refer anomalies as positive (+) samples, and normal data points as negative (-) samples.
3.1 Positive-Unlabeled Autoencoder
Let $p_{\mathcal{N}}$ be the normal data distribution, $p_{\mathcal{A}}$ be the seen anomaly distribution, and $p_{\mathcal{U}}$ be the unlabeled data distribution. We assume that the datasets $\mathcal{U}$ and $\mathcal{A}$ are drawn from $p_{\mathcal{U}}$ and $p_{\mathcal{A}}$, respectively. We also assume that $p_{\mathcal{U}}$ can be rewritten as follows:
$p_{\mathcal{U}}(\mathbf{x})=\alpha p_{\mathcal{A}}(\mathbf{x})+(1-\alpha)p_{\mathcal{N}}(\mathbf{x}),$ (6)
where $\alpha\in[0,1]$ represents the probability of anomaly occurrence in the unlabeled data. Hence, $p_{\mathcal{N}}$ can be rewritten as follows:
$(1-\alpha)p_{\mathcal{N}}(\mathbf{x})=p_{\mathcal{U}}(\mathbf{x})-\alpha p_{\mathcal{A}}(\mathbf{x}).$ (7)
Although $\alpha$ is the hyperparameter and is assumed to be known throughout this paper, it can be estimated from the datasets $\mathcal{U}$ and $\mathcal{A}$ in conventional PU learning approaches [29, 44, 17, 6].
If we have access to the normal data distribution $p_{\mathcal{N}}$, we can train the AE by minimizing the following ideal objective function:
$\mathcal{L}_{\mathrm{PN}}(\theta)=\alpha\mathbb{E}_{p_{\mathcal{A}}}[\ell_{\mathrm{BCE}}(\mathbf{x},1;\theta)]+(1-\alpha)\mathbb{E}_{p_{\mathcal{N}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)],$ (8)
where $\mathbb{E}[\cdot]$ is the expectation. Since we cannot access $p_{\mathcal{N}}$ in practice, we have to approximate the second term in Eq. (8). According to Eq. (7), this can be rewritten as follows:
$(1-\alpha)\mathbb{E}_{p_{\mathcal{N}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)]=\mathbb{E}_{p_{\mathcal{U}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)]-\alpha\mathbb{E}_{p_{\mathcal{A}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)].$ (9)
Hence, using the seen anomaly distribution $p_{\mathcal{A}}$ and the unlabeled data distribution $p_{\mathcal{U}}$, $\mathcal{L}_{\mathrm{PN}}(\theta)$ can be rewritten as follows:
$\mathcal{L}_{\mathrm{PN}}(\theta)=\alpha\mathbb{E}_{p_{\mathcal{A}}}[\ell_{\mathrm{BCE}}(\mathbf{x},1;\theta)]+\mathbb{E}_{p_{\mathcal{U}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)]-\alpha\mathbb{E}_{p_{\mathcal{A}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)].$ (10)
With the datasets $\mathcal{U}$ and $\mathcal{A}$, we can approximate $\mathcal{L}_{\mathrm{PN}}(\theta)$ by the Monte Carlo approximation as follows:
$\mathcal{L}_{\mathrm{PN}}(\theta)\simeq\alpha\underbrace{\frac{1}{M}\sum_{m=1}^{M}\ell_{\mathrm{BCE}}(\tilde{\mathbf{x}}_{m},1;\theta)}_{\mathcal{L}_{\mathcal{A}}^{+}(\theta)}+\underbrace{\frac{1}{N}\sum_{n=1}^{N}\ell_{\mathrm{BCE}}(\mathbf{x}_{n},0;\theta)}_{\mathcal{L}_{\mathcal{U}}^{-}(\theta)}-\alpha\underbrace{\frac{1}{M}\sum_{m=1}^{M}\ell_{\mathrm{BCE}}(\tilde{\mathbf{x}}_{m},0;\theta)}_{\mathcal{L}_{\mathcal{A}}^{-}(\theta)}.$ (11)
In this equation, the sum of the second and third terms is the approximation of the anomaly scores for normal data:
$(1-\alpha)\mathbb{E}_{p_{\mathcal{N}}}[\ell_{\mathrm{BCE}}(\mathbf{x},0;\theta)]\simeq\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta).$ (12)
The left-hand side in Eq. (12) is always greater than or equal to zero, but the right-hand side can be negative. In experiments, it often converges towards negative infinity, resulting in the meaningless solution. To avoid this, according to [22], our training objective function to be minimized ensures that $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ is not negative as follows:
$\mathcal{L}_{\mathrm{PUAE}}(\theta)=\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\max\left\{0,\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\right\}.$ (13)
Algorithm 1 Positive-Unlabeled Autoencoder
1:Unlabeled and anomaly datasets $(\mathcal{U},\mathcal{A})$, mini-batch size $K$, hyperparameter $\alpha\in[0,1]$
2:Model parameter $\theta$
3:while not converged do
4:     Sample mini-batch $\mathcal{B}$ from datasets $(\mathcal{U},\mathcal{A})$
5:     Compute $\mathcal{L}_{\mathcal{A}}^{+}(\theta)$, $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$, and $\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ in Eq. (11) with $\mathcal{B}$
6:     if $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\geq 0$ then
7:         Set the gradient $\nabla_{\theta}(\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta))$
8:     else
9:         Set the gradient $\nabla_{\theta}-(\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta))$
10:     end if
11:     Update $\mathbf{\theta}$ with the gradient
12:end while
We can optimize this training objective function by using the stochastic gradient descent (SGD) such as Adam [19]. In practice, according to [22], when $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)<0$, we maximize it to be greater than or equal to zero. We call our proposed method the positive-unlabeled autoencoder (PUAE). Algorithm 1 shows the pseudo code of the PUAE, where $K$ is the mini-batch size for the SGD. Note that our approach can be easily extended to positive-negative-unlabeled (PNU) learning [49], where we can also use a small amount of normal data.
3.2 Application to Other Anomaly Detectors
The PUAE is based on the loss function of the AE defined in Eq. (1), and can be applied to other anomaly detectors by substituting their loss function into Eq. (1). This application requires that their loss function be non-negative and differentiable. A lot of approaches satisfy this requirement, such as the denoising AE (DAE) [55] and deep support vector data description (DeepSVDD) [46].
One example is the DAE, which is a variant of the AE, and has also been successfully applied to anomaly detection [50]. The DAE aims to reconstruct original data points from noisy input data points. Its loss function for each data point is defined as follows:
$\ell(\mathbf{x};\theta)=\left\|D_{\theta}(E_{\theta}(\mathbf{x}+\epsilon))-\mathbf{x}\right\|,$ (14)
where $\epsilon$ is a noise from an isotropic Gaussian distribution. Our approach can be applied to the DAE by substituting this into Eq. (1).
The other example is the DeepSVDD, which aims to pull the representation of the normal data towards the predefined center, and push those of the anomaly data away from it. Let $f_{\theta}(\mathbf{x})$ be the feature extractor, like the encoder in the AE. The loss function for each data point of the DeepSVDD is defined as follows:
$\ell(\mathbf{x};\theta)=\left\|f_{\theta}(\mathbf{x})-\mathbf{c}\right\|^{2},$ (15)
where $\mathbf{c}\neq\mathbf{0}$ is the predefined center vector. Our approach can be applied to the DeepSVDD by substituting this into Eq. (1).
4 Related Work
4.1 Unsupervised Anomaly Detection
Numerous unsupervised anomaly detectors have been presented, ranging from shallow approaches such as the one-class support vector machine (OCSVM) [54] and the isolation forest (IF) [27] to deep approaches such as the AE [16] and the DeepSVDD [46]. In addition, generative models such as the variational autoencoder [20, 21] and the generative adversarial nets [12] are also used for anomaly detection [5, 52, 45, 40, 58, 14, 60]. Although they are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies. For example, generative models may fail to detect anomalies that are obvious to the human eye [32].
Furthermore, these approaches assume that unlabeled data are mostly normal. However, they are contaminated with anomalies in practice, degrading the detection performance. Several unsupervised approaches have been presented to handle such contaminated unlabeled data [63, 43, 53]. Among them, the latent outlier exposure (LOE) [43] is one of the most representative methods. The LOE introduces the label for each data point as the latent variable, and alternates between inferring the latent label and optimizing the parameter of the base anomaly detector. Compared to these unsupervised approaches, our approach can achieve high anomaly detection performance by using the unlabeled and anomaly data, even if the unlabeled data are contaminated with anomalies. As the base anomaly detector for our approach, the AE and the DeepSVDD can be used as described in Section 3. In addition, our approach can also be applied to the LOE by substituting its objective function into $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$ in Eq. (11).
4.2 Semi-supervised Anomaly Detection
Several semi-supervised anomaly detectors, aiming to improve the detection performance using anomaly data in addition to unlabeled data, have been presented [15, 59, 47]. Among them, the DeepSAD [47] is one of the most representative methods. The DeepSAD trains the DeepSVDD model to minimize Eq. (15) for the unlabeled data, and to maximize it for the anomaly data. Compared to these semi-supervised approaches, our approach can handle the unlabeled data that are contaminated with anomalies, as described in Section 3.1.
To handle contaminated unlabeled data, several semi-supervised approaches have been presented, including PU learning approaches [62, 18, 61, 37, 25, 41]. Among them, the semi-supervised outlier exposure with a limited labeling budget (SOEL) [25] is the current state-of-the-art approach. The SOEL is a semi-supervised extension of the LOE [43], and presents the query strategy for the LOE, deciding which data should be labeled. We compare our approach with the SOEL in Section 5, and confirm the effectiveness of our approach.
4.3 Positive-Unlabeled Learning
A lot of PU learning approaches have been presented in the context of binary classification [10, 9, 8, 22, 2, 31]. Among them, our approach is based on the unbiased PU learning [9, 8, 22]. The empirical risk estimators in Eq. (11) is unbiased and consistent with respect to all popular loss function. This means for fixed $\theta$, Eq. (11), which is the approximation of Eq. (8), converges to Eq. (8) as the dataset sizes $N,M\rightarrow\infty$. It is known that if the model is linear with respect to $\theta$, a particular loss function will result in convex optimization, and the globally optimal solution can be obtained [33, 38, 35]. Despite this ideal property, when the model is complex such as neural networks, Eq. (11) can be negative, potentially leading to the meaningless solution. To avoid this, according to [22], we use the max operator to prevent it from being negative.
Compared to conventional PU learning that is based on the binary classifier, our approach is based on the anomaly detector such as the AE. Although conventional PU learning cannot detect unseen anomalies since its decision boundary is between normal data points and seen anomalies, our approach can detect both seen and unseen anomalies.
5 Experiments
5.1 Data
We used following image datasets: MNIST [51], FashionMNIST [57], SVHN [34], and CIFAR10 [23]. MNIST is the handwritten digits, FashionMNIST is the fashion product images, SVHN is the house number digits, and CIFAR10 is the animal and vehicle images. We converted SVHN and CIFAR10 into gray-scale, and resized all datasets to $32\times 32$ resolution.
These datasets consist of 10 class images. For MNIST and SVHN, we used the digits as class indices. For FashionMNIST, we indexed labels as: {T-shirt/top: 0, Trouser: 1, Pullover: 2, Dress: 3, Coat: 4, Sandal: 5, Shirt: 6, Sneaker: 7, Bag: 8, and Ankle boot: 9}. For CIFAR10, we indexed labels as: {airplane: 0, automobile: 1, bird: 2, cat: 3, deer: 4, dog: 5, frog: 6, horse: 7, ship: 8, and truck: 9}.
We extend the experiments in [46] to semi-supervised anomaly detection with contaminated unlabeled data. Of the 10 classes, we used one class as normal, another class as unseen anomaly, and the remaining classes as seen anomaly. For example with MNIST, if we use the digit 1 as normal and the digit 0 as unseen anomaly, seen anomaly corresponds to the digits 2, 3, 4, 5, 6, 7, 8, and 9. For all datasets, we used class 0 as unseen anomaly, and select one normal class from the remaining 9 classes. The training dataset consists of 5,000 samples, of which 4,500 samples are unlabeled normal data points, 250 samples are labeled seen anomalies, and 250 samples are unlabeled seen anomalies. That is, the unlabeled data points in this dataset are contaminated with seen anomalies. We used 10% of the training dataset as the validation dataset. The test dataset consists of 2,000 samples, about half of which are normal and the rest are anomalies, including both seen and unseen anomalies. More precisely, the normal data points are sampled from the test dataset of the normal class up to 1,000 samples, and the rest are sampled from the other classes. We provide the example of the dataset in the case of MNIST in Appendix A.
5.2 Methods
We applied our approach to the AE [16] and the DeepSVDD [46]. We refer to the former as PUAE, and the latter as PUSVDD. We compared our PUAE and PUSVDD with the following unsupervised and semi-supervised approaches.
Unsupervised approaches: We used the IF [27] as the shallow approach, and used the AE [16] and the DeepSVDD [46] as the deep approaches. We also used the LOE [43], which is robust to the anomalies in the unlabeled data. We chose the AE as the base detector for the LOE.
Semi-supervised approaches: We used the ABC [59], the DeepSAD [47], and the SOEL [25] that is a semi-supervised extension of the LOE. We chose the AE as the base detector for the SOEL. We also used the PU learning binary classifier (PU) [22] for reference.
5.3 Setup
First, we outline the setups for deep approaches, which include all approaches except the IF. We used two-hidden-layer neural networks (1,000 units per layer) for the encoder and the decoder in the AE, the feature extractor in the DeepSVDD, and the binary classifier in the PU. We used the ReLU as the activations for the above neural networks. For the AE-based and DeepSVDD-based approaches, we set the dimension of the latent variable to 40. For the DeepSVDD-based approaches, we used no bias terms in each layer, pretrained these feature extractors as the AE, and set the center $\mathbf{c}$ in Eq. (15) to the mean of the outputs of the encoder. We trained all methods by using Adam [19] with a mini-batch size of 128. We set the learning rate to $10^{-4}$, set the maximum number of epochs to 200, used the weight decay [13] with $10^{-3}$, and used early-stopping [13] based on the validation dataset. We set the hyperparameter $\alpha=0.1$ for the PUAE, the PUSVDD, the PU, the LOE, and the SOEL, which represents the probability of anomaly occurrence.
We next outline the setup for the IF. We used the implementation provided by scikit-learn [39]. The hyperparameters were left as default in our experiments.
For unsupervised approaches, we trained them only using the unlabeled data in the training dataset. Note that we did NOT use the anomaly data since unsupervised approaches cannot effectively use them. For semi-supervised approaches, we trained them using the unlabeled and anomaly data.
To measure the anomaly detection performance, we calculated the AUROC scores for all datasets. We ran all experiments eight times while changing the random seeds.
The machine specifications used in the experiments are as follows: the CPU is Intel Xeon Platinum 8360Y, the memory size is 512GB, and the GPU is NVIDIA A100 SXM4.
5.4 Results
Table 1: Comparison of anomaly detection performance on image datasets. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value.
MNIST FashionMNIST SVHN CIFAR10
IF [27] 0.829 $\pm$ 0.089 0.911 $\pm$ 0.060 0.514 $\pm$ 0.013 0.561 $\pm$ 0.097
AE [16] 0.894 $\pm$ 0.056 0.823 $\pm$ 0.094 0.598 $\pm$ 0.027 0.532 $\pm$ 0.127
DeepSVDD [46] 0.763 $\pm$ 0.060 0.667 $\pm$ 0.070 0.513 $\pm$ 0.014 0.556 $\pm$ 0.032
LOE [43] 0.905 $\pm$ 0.050 0.860 $\pm$ 0.101 0.594 $\pm$ 0.028 0.533 $\pm$ 0.125
ABC [59] 0.897 $\pm$ 0.055 0.826 $\pm$ 0.093 0.599 $\pm$ 0.027 0.533 $\pm$ 0.127
DeepSAD [47] 0.795 $\pm$ 0.054 0.693 $\pm$ 0.072 0.514 $\pm$ 0.016 0.560 $\pm$ 0.033
SOEL [25] 0.944 $\pm$ 0.034 0.896 $\pm$ 0.075 0.601 $\pm$ 0.032 0.535 $\pm$ 0.123
PU [22] 0.943 $\pm$ 0.071 0.811 $\pm$ 0.180 0.500 $\pm$ 0.014 0.488 $\pm$ 0.066
PUAE 0.971 $\pm$ 0.019 0.940 $\pm$ 0.047 0.608 $\pm$ 0.030 0.554 $\pm$ 0.115
PUSVDD 0.922 $\pm$ 0.029 0.884 $\pm$ 0.060 0.538 $\pm$ 0.030 0.575 $\pm$ 0.037
Table 1 compares the anomaly detection performance on each image dataset. First, we focus on the unsupervised approaches. The IF showed the high detection performance on FashionMNIST and CIFAR10. This indicates that the tree structure is somewhat effective on the contaminated datasets. The deep approaches such as the AE and the DeepSVDD are more easily influenced by the contaminated datasets due to their flexibility. As a result, they performed worse than our PUAE and PUSVDD in all cases. The DeepSVDD performed worse than the AE except for CIFAR10. This indicates that the DeepSVDD is more easily influenced by the contaminated datasets than the AE. One possible reason is that the training objective function of the DeepSVDD (Eq. (15)) is much simpler than that of the AE (Eq. (1)). Although the LOE does not use the anomaly data, it can mitigate the effect of the contaminated datasets to some extent. In particular, it significantly improved the detection performance in FashionMNIST compared to the AE, which is the base detector of the LOE.
Next, we focus on semi-supervised approaches. The ABC and the DeepSAD improved the detection performance compared to the AE and the DeepSVDD, respectively. This indicates that we can improve the detection performance by using the anomaly data. However, their improvements were marginal because the unlabeled data are contaminated with anomalies. The SOEL, which can handle such unlabeled data, achieved the second best performance compared to our approaches. These results indicate the effectiveness of our approaches in handling the contaminated unlabeled data.
Finally, we focus on our approaches. In almost all cases, our PUAE and PUSVDD performed better than the AE and the ABC, and the DeepSVDD and the DeepSAD, respectively. Especially, in MNIST, FashionMNIST, and SVHN, the PUAE achieved the best performance among all detectors. This shows the effectiveness of our approach, which is based on PU learning and the anomaly detector. Although the PUAE performed poorly in CIFAR10, the PUSVDD achieved the best performance in CIFAR10. This indicates that we can achieve higher detection performance by selecting the base anomaly detector that is better suited to the target dataset.
In addition, we focus on the difference between our approaches and the PU. Although the PU achieved high performance on MNIST and FashionMNIST, its performance was worse than the PUAE in all datasets. The reason is as follows. Since the PU is for binary classification, it sets the decision boundary between normal data points and seen anomalies. This prevents us from detecting unseen anomalies. On the other hand, our approach can detect unseen anomalies since it can model normal data points by the anomaly detector. To support this claim, we provide the anomaly detection performance for seen and unseen anomalies on each dataset in Appendix B.
5.5 Hyperparameter Sensitivity
(a) MNIST
(b) FashionMNIST
(c) SVHN
(d) CIFAR10
Figure 2: Relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUAE on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies. We plotted the AUROC of the ABC by a dashed line for comparison. The semi-transparent area represents standard deviations. The anomaly detection performance is improved by tuning $\alpha$.
Our approach has the hyperparameter $\alpha$, which represents the probability of anomaly occurrence. In the above experiments, we set it to $\alpha=0.1$ since the 10% of the training dataset is seen anomalies. Although $\alpha$ can be estimated from the unlabeled and anomaly training data in conventional PU learning approaches [29, 44, 17, 6], its nature in our approach has not yet been clarified.
Finally, we evaluate the sensitivity of $\alpha$ in our approach. Figure 2 shows the relationship between the anomaly detection performance and the hyperparameter $\alpha$ of the PUAE on each dataset. We used class 0 as unseen anomaly, class 1 as normal, and the remaining classes as seen anomalies on each dataset.
It can be seen that the performance is improved by tuning $\alpha$. Note that the performance is better when $\alpha$ is set to a larger value than $0.1$, which is the true probability of anomaly occurrence. The possible reasons for this are as follows. In the objective function of our approach (Eq. (13)), the larger $\alpha$, the greater the effect of maximizing the anomaly scores for the anomaly data. In the case of conventional PU learning for binary classification, if $\alpha$ is set unnecessarily large, the decision boundary of the classifier will be biased. On the other hand, since the PUAE is based on the autoencoder, it is possible to maintain the ability to reconstruct normal data points even if $\alpha$ becomes large to some extent. Therefore, for improving the anomaly detection performance, it is better to increase $\alpha$ so that the PUAE fails to reconstruct anomalies. At present, we should determine $\alpha$ based on the validation dataset. The analysis regarding the optimal $\alpha$ is part of our future work.
The above experiments and discussion show that our approach is effective on various datasets, and that we can further improve its performance by selecting the appropriate base detector and tuning the hyperparameter.
6 Conclusion and Limitations
Although unlabeled data are assumed to be mostly normal in anomaly detection, they are often contaminated with anomalies in practice, which prevents us from improving the detection performance. To solve this, we propose the positive-unlabeled autoencoder, the semi-supervised anomaly detection approach based on PU learning and the anomaly detector such as the AE and the DeepSVDD. Our approach enables us to approximate the anomaly scores for normal data with the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the anomaly detector to minimize the anomaly scores for normal data, and to maximize those for the anomaly data. Our approach achieves better detection performance than existing approaches on various datasets.
Our approach has the following two limitations. First, our approach assume that the unlabeled data do not contain unseen anomalies. We can handle such unseen anomalies by using the robust anomaly detector for our approach, such as the LOE. Second, our approach has a hyperparameter that represents the probability of anomaly occurrence. Although this hyperparameter can be estimated by conventional PU learning approaches, we can further improve the detection performance by tuning it, as discussed in Section 5.5. Analyzing this hyperparameter is our important future work.
Appendix A Dataset Example
(a) Training (unlabeled)
(b) Training (anomaly)
(c) Test (normal)
(d) Test (anomaly)
Figure 3: The example of the dataset in the case of MNIST.
Figure 3 shows the example of the dataset in the case of MNIST. In this example, the digit 1 is normal, the digit 0 is unseen anomaly, and the digits 2, 3, 4, 5, 6, 7, 8, and 9 are seen anomalies. (a) The unlabeled data points in the training dataset are contaminated with seen anomalies. (b) The anomaly data points in the training dataset contain seen anomalies but not unseen anomalies. (c) The normal data points in the test dataset are not contaminated with anomalies. (d) The anomaly data points in the test dataset contain both seen and unseen anomalies.
Appendix B Additional Results
Table 2: Comparison of anomaly detection performance for seen anomalies.
MNIST FashionMNIST SVHN CIFAR10
IF [27] 0.814 $\pm$ 0.095 0.913 $\pm$ 0.056 0.514 $\pm$ 0.015 0.566 $\pm$ 0.098
AE [16] 0.882 $\pm$ 0.062 0.822 $\pm$ 0.097 0.597 $\pm$ 0.024 0.545 $\pm$ 0.125
DeepSVDD [46] 0.758 $\pm$ 0.061 0.661 $\pm$ 0.071 0.514 $\pm$ 0.016 0.556 $\pm$ 0.033
LOE [43] 0.893 $\pm$ 0.056 0.861 $\pm$ 0.105 0.593 $\pm$ 0.026 0.547 $\pm$ 0.125
ABC [59] 0.885 $\pm$ 0.061 0.825 $\pm$ 0.097 0.598 $\pm$ 0.024 0.546 $\pm$ 0.125
DeepSAD [47] 0.793 $\pm$ 0.055 0.690 $\pm$ 0.072 0.516 $\pm$ 0.016 0.559 $\pm$ 0.036
SOEL [25] 0.937 $\pm$ 0.039 0.899 $\pm$ 0.077 0.600 $\pm$ 0.029 0.549 $\pm$ 0.123
PU [22] 0.951 $\pm$ 0.063 0.817 $\pm$ 0.182 0.503 $\pm$ 0.013 0.509 $\pm$ 0.069
PUAE 0.970 $\pm$ 0.021 0.943 $\pm$ 0.041 0.608 $\pm$ 0.027 0.568 $\pm$ 0.115
PUSVDD 0.929 $\pm$ 0.028 0.891 $\pm$ 0.054 0.540 $\pm$ 0.032 0.576 $\pm$ 0.039
Table 3: Comparison of anomaly detection performance for unseen anomalies.
MNIST FashionMNIST SVHN CIFAR10
IF [27] 0.954 $\pm$ 0.033 0.917 $\pm$ 0.112 0.495 $\pm$ 0.012 0.533 $\pm$ 0.077
AE [16] 0.996 $\pm$ 0.003 0.847 $\pm$ 0.113 0.607 $\pm$ 0.037 0.431 $\pm$ 0.104
DeepSVDD [46] 0.806 $\pm$ 0.065 0.712 $\pm$ 0.081 0.515 $\pm$ 0.015 0.558 $\pm$ 0.031
LOE [43] 0.996 $\pm$ 0.003 0.873 $\pm$ 0.119 0.606 $\pm$ 0.036 0.429 $\pm$ 0.104
ABC [59] 0.996 $\pm$ 0.004 0.848 $\pm$ 0.113 0.607 $\pm$ 0.038 0.432 $\pm$ 0.104
DeepSAD [47] 0.815 $\pm$ 0.065 0.723 $\pm$ 0.084 0.515 $\pm$ 0.014 0.560 $\pm$ 0.030
SOEL [25] 0.996 $\pm$ 0.003 0.887 $\pm$ 0.119 0.608 $\pm$ 0.041 0.432 $\pm$ 0.102
PU [22] 0.877 $\pm$ 0.181 0.774 $\pm$ 0.223 0.491 $\pm$ 0.010 0.316 $\pm$ 0.047
PUAE 0.978 $\pm$ 0.015 0.915 $\pm$ 0.118 0.612 $\pm$ 0.041 0.452 $\pm$ 0.097
PUSVDD 0.868 $\pm$ 0.074 0.846 $\pm$ 0.125 0.531 $\pm$ 0.027 0.565 $\pm$ 0.036
As the additional experiments, we evaluate the anomaly detection performance for seen and unseen anomalies on each dataset. The setups are the same as those described in Section 5. For the test data, we sampled normal data points, seen anomalies, and unseen anomalies from the test dataset up to 1,000 samples each.
Tables 2 and 3 show the anomaly detection performance for seen and unseen anomalies, respectively. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value.
For seen anomalies, the PUAE achieved the best performance among all detectors. This shows the effectiveness of our approach, which is robust to the contaminated unlabeled data according to PU learning. For unseen anomalies, the PUAE achieved the best performance on FashionMNIST and SVHN, and the PUSVDD achieved the best performance on CIFAR10. This indicates that we may be able to improve the detection performance for unseen anomalies by using seen anomalies.
These results also show the difference between the conventional PU learning and our approach. The PU achieved the poor detection performance for unseen anomalies. This is because that it sets the decision boundary between normal data points and seen anomalies. On the other hand, our approach can detect unseen anomalies since it is based on the anomaly detector.